what is the length of the diagonal of a regular cube with a surface area of 54 units squared?
Answer by mananth(16949)

surface area of cube = 54
let the side be x
x^2 is the area of one face of the cube.
there are 6 faces.
6x^2 = 54
x^2= 54/6
x^2=9
x=3 units
apply pythagoras theorem to find the diagonal.
3^2 +3^2 = diagonal^2
9+9= diagonal ^2
diagonal = sqrt18